OpenStudy (anonymous):

Solve 7^(x - 2) = 12^x

OpenStudy (anonymous):

Gotta use log...

OpenStudy (anonymous):

i know that much, but the thing is I got the answer, x = 3.28, and not the right answer of -7.22, if you could just give me the step by step process that would be great

OpenStudy (anonymous):

Okay:)

OpenStudy (anonymous):

ok so here's what I did, just instruct me on what i did wrong

OpenStudy (anonymous):

7^(x-2)=12^(x) ln(7^(x-2))=ln(12^(x)) xln(7)-2ln(7)=ln(12^(x)) xln(7)-2ln(7)-xln(12)=0 x(ln(7)-ln(12))=2ln(7) x=(2ln(7))/(ln(7)-ln(12)) x=3.8918

OpenStudy (anonymous):

\[(x-2)(\log(7)) =(x) \log(12) \]

OpenStudy (anonymous):

\[(x-2) (.8450984) = x(1.079181246)\]

OpenStudy (anonymous):

Uh hu...The sooner you turn it into a decimal the faster it gets farther from an exact answer. Just sayin. I think you did it correctly just rounding got the best of you!

OpenStudy (anonymous):

I"m not really rounding, I'm just putting out what the calc it putting out

OpenStudy (anonymous):

it's worked for all the other ones

OpenStudy (anonymous):

Well the calculator is rounding lol.

OpenStudy (anonymous):

yeah i guess the calc rounds

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Thanks though! I get it now

OpenStudy (anonymous):

Good:)